# Beyond Community Notes: A Framework for Understanding and Building Crowdsourced Context Systems

### v1

@@ 5.2. Design Feature Identification
The researchers independently read through the articles one at a time and extracted design details into a spreadsheet, which resulted in a set of “design features” for each platform. After independently reviewing the articles for a platform, the researchers convened and performed a round of affinity diagramming (Beyer and Holtzblatt, 1999), in which they wrote the design features on sticky notes and placed them on a whiteboard. They placed “similar” design features together, and through this inductive method identified “feature groups”. The researchers then performed thematic analysis (Braun and Clarke, 2006) to identify the themes of the groupings and came up with provisional group names. They repeated this process with the articles for each platform, reconvening to do another round of affinity diagramming after each round of feature extraction. These design features and feature groups were further synthesized into the aspects, dimensions, and options of the design space, which we present below in Section 7.
@@
6. What Does the Research Community Know About CCS?
Study Focus Primary Method Sample Paper
System Impact (n=10) Data Analysis: X Community Notes Data Diffusion of Community Fact-Checked Misinformation on Twitter (Drolsbach and Pröllochs (2023))
Contributor Behavior (n=9) Data Analysis: X Community Notes Data Birds of a feather don’t fact-check each other: Partisanship and the evaluation of news in Twitter’s Birdwatch crowdsourced fact-checking program (Allen et al. (2022))
User’s Perceptions (n=6) User studies Checking the Fact-Checkers: The Role of Source Type, Perceived Credibility, and Individual Differences in Fact-Checking Effectiveness (Liu et al. (2023))
System Design (n=5) User studies, design specs, algorithm analysis Supernotes: Driving Consensus in Crowd-Sourced Fact-Checking (De et al. (2025))
Table 1. Summary of the existing research found in our literature review.:
@@ 6. What Does the Research Community Know About CCS?
Our systematic literature review identified n=30 papers about Crowdsourced Context Systems, summarized in Table 1. These papers fall into several high-level categories: studies of the system’s impact, contributor behavior, users’ perceptions of crowdsourced context, and alternative system designs. Some studies contained elements of multiple categories, in which case we grouped them based on their most significant contribution.
@@
Many of the published CCS research studies focused on the CCS’s impact on the platform where it is implemented (n=10 papers). These studies all performed analysis of data from X’s Community Notes. These studies asked both causal questions, such as about the effect that displaying a note has on post diffusion (Drolsbach and Pröllochs, 2023; Chuai et al., 2024b), as well as descriptive questions about the performance of X’s system in several specific use cases (Chujyo et al., 2025; Bouchaud and Ramaciotti, 2025). The studies found strong evidence that CCS can both reduce the spread (Drolsbach and Pröllochs, 2023) and believability (Wojcik et al., 2022) of misleading posts. The research also argued that the bridging-based algorithm in X’s Community Notes system has demonstrated it can identify high-quality crowdsourced notes that are likely to be perceived as helpful by users across an ideological spectrum (Bouchaud and Ramaciotti, 2025). However, the studies also showed that this bridging approach takes a significant amount of time to identify “helpful” notes for misleading posts. As a result, there is often a significant delay before notes are displayed to users (Renault et al., 2024; Chuai et al., 2024a, b). Additionally, studies have shown that the vast majority of suggested notes ($\sim 90\%$) are never published due to either a lack of ratings or consensus  (Chujyo et al., 2025; Wirtschafter and Majumder, 2023).
@@
Another common type of research study (n=9) examined CCS contributor behavior. While one of these studies used interviews as a data source (Yoon et al., 2025), all of the others relied on the open data from X’s Community Notes. The studies asked descriptive questions about how contributors to the system behaved, such as what types of posts (Pilarski et al., 2023; Saeed et al., 2022) and accounts (Allen et al., 2022) contributors tend to leave notes on, and what type of content contributors tend to include in notes (Jones et al., 2022; Kuuse et al., 2025). Two key findings from these studies are that 1) the crowd exerts political bias in the posts that they select to write notes for and the ratings that they assign to notes (Allen et al., 2022); and 2) the crowd behaves differently than professional fact-checkers (Pilarski et al., 2023), yet often links to them as sources (Borenstein et al., 2025).
A third group of studies sought to understand user’s perceptions of CCS (n=6). These studies used experiments (Liu et al., 2023), interviews (Shusas and Forte, 2024), and data analysis (Hameleers, 2024) of X’s data to answer descriptive questions about how individuals perceived the credibility and helpfulness of context provided by CCS. These studies have demonstrated that certain types of notes are perceived as more helpful than others, such as those that support their claims with links to unbiased (i.e., assessed as politically neutral by a third-party organization [^9] https://mediabiasfactcheck.com/) external sources (Solovev and Pröllochs, 2025). There is also evidence that users perceive notes from CCS as just as, or even more, credible than those from professional fact checkers (Liu et al., 2023; Shusas and Forte, 2024).
@@
The final group of studies were design studies that either proposed or built a variation of a CCS (n=5). These papers focused on ways to improve the quantity (Li et al., 2025) and quality (Mohammadi and Yasseri, 2025) of notes, as well as techniques to improve the performance of the note curation algorithm (Mujumdar and Kumar, 2022). The bridging-based algorithm for selecting “helpful” notes was introduced in one of these papers (Wojcik et al., 2022), as a solution to problems of bias (Allen et al., 2022) in a majority-rule note selection system. Recent papers have proposed ways of incorporating AI into the note-writing process and found that LLMs have the potential to generate helpful notes (De et al., 2025; Mohammadi and Yasseri, 2025). While these studies begin to imagine alternative system designs, much is still unexplored. As we show next, the potential design space is much wider than what is covered in these studies.
@@ 7.1. Participation
Dimension Implementing Platforms
@@
All users. Meta, Twitter/X
Only users from a specific geo/market. TikTok, YouTube
@@
Users. TikTok, Meta, Twitter/X, YouTube
@@
Users. TikTok, Meta, Twitter/X
Platform employees/contractors. YouTube
@@
Nothing: users cannot be context contributors. YouTube
Apply. TikTok, Meta, Twitter/X
Receive an invite from the platform. YouTube
Be admitted by the platform. TikTok, Meta, Twitter/X
Meet a quality threshold of past performance. Twitter/X
@@
Apply. TikTok, Meta, Twitter/X, YouTube
Receive an invite from the platform. YouTube
Be admitted by the platform. TikTok, Meta, Twitter/X, YouTube
@@
User-age requirements. TikTok, Meta
Account-age requirements. TikTok, Meta, Twitter/X, YouTube
Lack of policy-violations. TikTok, Meta, Twitter/X, YouTube
Geographic requirements. TikTok, Meta, Twitter/X
Language requirements. Meta, YouTube
Account authentication requirements. TikTok, Meta, Twitter/X
@@
Yes: platform has discretion over who they admit. TikTok, Meta, Twitter/X, YouTube
@@
Paid by the platform. YouTube
Pseudonymous reputation. Twitter/X
@@
Contributors can add notes to posts they organically come across while using the platform. TikTok, Meta, Twitter/X, YouTube
Contributors have feeds of posts that need notes, and notes that need ratings. Meta, Twitter/X
Platform sends notifications to contributors requesting review of specific posts and notes. Twitter/X
Users can request notes on posts or additional review on published notes. Meta, Twitter/X, YouTube
@@ 7.2. Inputs
Dimension Implementing Platforms
@@
Free-text inputs TikTok, Meta, Twitter/X, YouTube
Ability to add source links. Meta, YouTube
Ability to add references to a portion of the post (e.g. timestamp in the video being noted). YouTube
@@
A fixed set of options (i.e., helpful, somewhat helpful, not helpful) to select from. TikTok, Meta, Twitter/X, YouTube
A free-text explanation field. YouTube
A fixed set of explanation options. Meta, Twitter/X
@@
All. Twitter/X
Posts by specific account types (public, verified, pages). TikTok, Meta
Posts of specific types of content (no ads, none that may feature minors). YouTube
@@
Must adhere to community guidelines. TikTok, Meta, Twitter/X, YouTube
Length restrictions. Meta, YouTube
Cannot be edited once submitted. Meta, Twitter/X
@@
Must adhere to community guidelines. YouTube
Length restrictions. YouTube
Cannot be edited once submitted. Twitter/X, YouTube
Cannot be deleted once submitted. Twitter/X, YouTube
@@
Published notes can be hidden by the crowd indefinitely. TikTok, Meta, YouTube
Note authors can edit or delete their notes. Meta, YouTube
@@
Published notes are ”locked” after a certain amount of time. Twitter/X
@@ 7.3. Curation
Dimension Implementing Platforms
@@
Algorithmic decision based on data from contributors. TikTok, Meta, Twitter/X, YouTube
@@
Platform can remove notes if they violate platform content policy. TikTok, Meta, Twitter/X, YouTube
@@
Rating contributor’s past platform behavior as a contributor. Twitter/X
@@
Appeal across the ideological spectrum (Bridging). TikTok, Meta, Twitter/X, YouTube
Other?
@@
Perhaps the most important design question is “What process determines which notes are shown alongside posts?” The existing systems that have recently grown in popularity all involve algorithmic decisions by bridging-based algorithms that, broadly speaking, identify notes that have been rated as “helpful” by contributors who have disagreed in their past ratings. Even though these bridging algorithms are currently popular, alternatives to this scheme are possible: for example, Twitter’s first implementation of Birdwatch used a majority voting algorithm. Alternatively, it could be possible to achieve bridging goals using contributor information other than their past ratings. Algorithmic approaches are also not the only option, as one could imagine a CCS that selects notes via editorial decisions by either the platform or contributors. Another possible approach is to select notes via digital juries (Fan and Zhang, 2020) of users, which could be selected either randomly or based on other criteria.
@@ 7.4. Presentation
Dimension Implementing Platforms
@@
Below the text of the original post. TikTok, Meta, Twitter/X, YouTube
@@
On both the original post and any re-shares of it. Twitter/X
@@
Original post authors. TikTok, Meta, Twitter/X, YouTube
Contributors who have written/rated the published note. Meta, YouTube
Every user that replied, liked, or reposted the original post. Meta, Twitter/X
@@
At most one. TikTok, Meta, Twitter/X, YouTube
@@
The author of the note is anonymous/omitted completely. TikTok, Meta, Twitter/X, YouTube
The author of the note is pseudonymous. Twitter/X
@@
Only whether or not it is “helpful”. TikTok, Meta, Twitter/X, YouTube
@@ 7.5. Platform Treatment
Dimension Implementing Platforms
@@
No different from any other posts. TikTok, Meta, Twitter/X, YouTube
@@
Limit positioning next to ads. Meta
@@
Replaces third-party fact checking program. Meta, Twitter/X
Complements platform-provided content moderation. TikTok, Meta, Twitter/X, YouTube
Notes are presented on posts alongside other “information reliability” indicators. TikTok, YouTube
@@ 7.6. Transparency
Dimension Implementing Platforms
@@
Open source. Twitter/X
Open data. Twitter/X
@@
Not public. TikTok, Meta, Twitter/X, YouTube
@@
Available to paying customers. Twitter/X
Not public. TikTok, Meta, YouTube
@@
Blog posts. TikTok, Meta, YouTube
Help articles/FAQs. Meta
Technical documentation. Twitter/X
@@
8. Normative Concerns
CCS are complex sociotechnical systems (Cherns, 1976) that integrate people, algorithms, and speech. These systems’ growing influence over the information shown on major social media platforms means that the design choices made in their implementation can have significant social consequences (Friedman and Nissenbaum, 1996; Flanagan et al., 2008). With such high stakes, it is crucial that a human-centered research agenda consider the implications of different design choices, in order to holistically evaluate the impacts of these novel systems.
The design space that we present in Section  7 highlights some of the crucial choices that are made when designing CCS. How are we to weigh the merits of different designs? One approach is to identify the primary normative concerns of such a system. We emphasize three interconnected normative concerns that should guide CCS design: user informedness, distribution of power, and fairness.
8.1. User Informedness
While social media serves many different purposes, CCS are specifically concerned with social media’s ability to inform the public. Such systems should thus be evaluated by how well they facilitate user informedness: for example, the degree to which they inform users about the accuracy of information, minimize the spread of misleading posts, or otherwise improve information quality on a platform. Our literature review surfaced ample evidence that the notes produced by CCS do have the power to increase user informedness, both by decreasing an individual’s belief in misleading information, and by decreasing the spread of this information (Wojcik et al., 2022; Renault et al., 2024; Slaughter et al., 2025; Chuai et al., 2024a, b; Drolsbach and Pröllochs, 2023). But while published notes have been shown to be effective at promoting understanding, we must also question how effective CCS are at producing these notes. Several of these same studies found that Twitter’s implementation of CCS produces notes too slowly to prevent the spread of the most viral content (Chuai et al., 2024b, a; Renault et al., 2024). There is also a significant gap in the research on the question of coverage, or how many of the misleading posts on a platform receive “helpful” notes, which is currently unanswered. The research suggests that there is a CCS design trade-off between speed/coverage and quality/helpfulness: X’s implementation has optimized for the latter, perhaps at the expense of the former. Whether this is the right decision should be up for debate.
For designers of CCS hoping to improve the system’s speed and coverage, there are several aspects from our design space that could be relevant. For example, within the curation aspect, designers could explore bridging algorithm changes that could lead to faster identification of “helpful” notes. Additionally, design changes within the participation and inputs aspects, such as those that would allow a larger contributor base or that would actively solicit notes on the most viral posts, could lead to a higher volume of notes and ratings, and thus quicker selection of helpful notes. When it comes to the difficulty of assessing the coverage of CCS, design changes within the transparency aspect could make this easier to study, perhaps by providing easier access to post engagement metrics.
8.2. Distribution of Power
One of the selling points of CCS is that they shift power over content moderation from the platform to the community. Giving users more of a say over the content that they see on their platform can make moderation decisions more legitimate and trusted (Shusas and Forte, 2024; Liu et al., 2023; Drolsbach et al., 2024), which is an admirable goal. Yet such claims require further scrutiny, and CCS should be evaluated to the extent that they actually empower users and do not merely allow platforms to shirk responsibility for misleading content. Evidence from our literature review suggests that bridging-based algorithms, by design, fail to produce notes for the most polarizing content (Bouchaud and Ramaciotti, 2025). For example, X’s implementation of the bridging algorithm requires polarized contributors to agree that a note is helpful in order to publish it. The editorial function of the bridging algorithm can thus limit the type and scope of contributions in two important ways: in terms of framing, the algorithm may prefer centrist points of view and content that could be agreeable to both sides of an issue; in terms of selection, agreement on a note may not occur for contentious topics, preventing its publication. Whether or not users at the ends of the ideological spectrum should be empowered to make moderation decisions is a valid question, but it is uncontroversial that the most polarizing content needs moderation, too. Research has found evidence of harmful content being handled poorly by the system, either because potential notes never reach helpful status (Chujyo et al., 2025), or because notes that do are not actually helpful for certain types of content, such as disinformation that is couched in humor (Matamoros-Fernández and Jude, 2025). Zooming out, the low publishing rate for suggested notes (around  10% on X (Wirtschafter and Majumder, 2023) and  6% on Meta [^10] https://x.com/guyro/status/1965800300557857053) gives us pause that the system is actually empowering these note authors. A related, underexplored area for research involves studying which note contributors are empowered by CCS. This research area is difficult to study holistically, as these systems are pseudonymous, but some work has begun to explore it qualitatively (Yoon et al., 2025) and via partnerships with platforms (Allen et al., 2022). More research is needed to understand if CCS empower users of certain demographics or political views over others.
Our design space offers several ways to further empower users via CCS. First and foremost is innovation in the participation aspect that would increase the set and diversity of contributors. The challenge for the platforms will be ensuing that quality remains as the participating contributor base increases, which is a tension that has been thoroughly studied in the crowdsourcing literature (Daniel et al., 2018), for example. Innovation in the inputs aspect may increase the quality of notes across the board, perhaps by giving note authors helpful feedback (Mohammadi and Yasseri, 2025), empowering more users to make successful contributions to the system. Another route to shift power to users is to disempower the platform. Though CCS are often described as ways to reduce platforms’ centrality in content moderation, platforms still exert power over the system in important ways, such as by deciding who is admitted as a contributor. Platforms could further guarantee their neutrality through changes to the transparency aspect, such as by making their contributor selection process public. But considering that X’s CCS implementation is the only even partially open system that we examined, a first step is for other platforms to follow X’s lead and open source their note curation algorithm and the crowdsourced data that they collect.
8.3. Fairness
CCS are often touted for their potential to implement content moderation in an unbiased way. Yet these systems are not inherently free from bias. Results from our literature review detail bias in contributor behavior, in terms of the types of accounts that they leave notes on, as well as how they rate suggested notes (Allen et al., 2022). Bridging algorithms have helped mitigate bias in the note curation process, with the caveats noted above, but additional bias can still be introduced—for example, in the note creation process, or in the selection of contributors. To give a concrete example: X’s bridging algorithm finds agreement among contributors, not the entire user base. If contributors’ beliefs are not representative of users’ beliefs then the algorithm’s outcomes may be a biased representation of the community, and not necessarily have broad appeal among users. There is some evidence that so far X’s implementation has not suffered from this problem, but it is worth considering. Also, accusations of bias in top-down moderation are often due to the opacity of such approaches and may be mitigated by transparency efforts. Only X’s implementation has provided any transparency, and yet the lack of transparency about the contributor base has made analyzing them for bias impossible. Finally, the approach of displaying context rather than removing posts is arguably fairer in that it lessens the damage from false positives, by empowering users to come to their own conclusions about content’s veracity (Peng and Grimmelmann, 2024). However, there is a trade off: leaving up questionable content, even with helpful context, can potentially do harm to those targeted by the content.
Several aspects from our design space can influence the fairness properties of a CCS. First off, all platforms have taken the approach that it is fairer not to remove misleading content. However designers may explore other options from the platform treatment dimension to identify certain cases where the platform should take action on content flagged by the crowd as misleading. Second, current implementations tend to display a single, definitive note, even if there may be several that are determined to be “helpful”. The presentation aspect offers alternatives that could further the aim of allowing users to come to their own conclusions, such as by showing multiple notes when there are multiple with helpful ratings. Additionally showing information about the note author could be helpful, as it could help users identify the credibility of the source. This may be at odds with contributor privacy, which current implementations have opted to protect. Finally, innovations in the curation aspect could explore other methods of selecting notes to display that may be regarded as fairer. Current implementations of the bridging algorithm use past ratings to identify contributors with diverse perspectives: perhaps it would be fairer to instead use in-app behavior? This could provide a fuller picture of a contributor’s ideological leanings. Encouraging more transparency would also give assurances of fairness, both in how contributors are selected, and how notes are published.

### v2

@@ 5.2. Design Feature Identification
The researchers independently read through the articles one at a time and extracted design details into a spreadsheet, which resulted in a set of design features for each platform. After independently reviewing the articles for a platform, the researchers convened and performed a round of affinity diagramming (beyer99), in which they wrote the design features on sticky notes and placed them on a whiteboard. They placed similar design features together, and through this inductive method identified feature groups. The researchers then performed thematic analysis (braunUsingThematicAnalysis2006) to identify the themes of the groupings and came up with provisional group names. They repeated this process with the articles for each platform, reconvening to do another round of affinity diagramming after each round of feature extraction. These design features and feature groups were further synthesized into the aspects, dimensions, and options of the design space, which we present below in Section 7.
@@
6. What Does the Research Community Know About Twitter-Style CCS?
Study Focus Primary Method Papers
System Impact (n=24) Data Analysis: X Community Notes Data (bouchaudAlgorithmicResolutionCrowdsourced2025b; renaultCollaborativelyAddingContext2024a; chuai25; truong2025communitynotesvulnerablerater; slaughter25; chujyo25; chuaiCommunitybasedFactcheckingReduces2024a; rollout21; kimDifferentialImpactIndividual2025e; diffusion21; chuai2025requestnoterequestfunction; matamoros25; razuvayevskaya2025timelinessconsensuscompositioncrowd; kangur2024checkscheckersexploringsource; DeepDiveXs; gao24; jama24; bobek2025communityfactchecksbreakfollower; borwankarDemocratizationMisinformationMonitoring2022a; Wirtschafter_Majumder_2023; wangEfficiencyCommunityBasedContent2024b; colemanLimitingFactorsEffectiveness2023; dekeulenaarTwitterDemotionCommunity2025; cham)
Contributor Behavior (n=16) Data Analysis: X Community Notes Data (allen22; borenstein2025; pilarskiCommunityNotesVs2023; pröllochs2021communitybasedfactcheckingtwittersbirdwatch; cikm22; kuuse25; jones22; solovevReferencesUnbiasedSources2025a; arjmandilari2025threatssustainabilitycommunitynotes; toyoda25; Phillips2025; renault25; fraxanetUnpackingPolarizationAntagonism2024c; martelPoliticalMotivesHelp2025; mohammadi2025birdwatchcommunitynotestwitter; yoon25)
System Design (n=10) User studies, design specs, algorithm analysis (mohammadiAIFeedbackEnhances2025a; wu2025crowdllmaugmentedcommunitynotes; wojcikBirdwatchCrowdWisdom2022b; hawkeye; franzmeyer2024hellofreshllmevaluationsstreams; liScalingHumanJudgment2025a; supernotes; zhang2025commenotessynthesizingorganiccomments; singh2025on; xing2025communitynotesdatasetexploringhelpfulness)
Users’ Perceptions (n=6) User studies (sharevski2025helps; liu23; shusas24; drolsbachCommunityNotesIncrease2024a; Kankham18062025; Hameleers18052024)
Table 1. The n=56 articles found in our literature review, grouped by study focus.:
@@ 6. What Does the Research Community Know About Twitter-Style CCS?
Our systematic literature review of research into Twitter-style CCS found n=56 papers and identified four high-level categories of study: studies of the system’s impact, contributor behavior, alternative system designs, and users’ perceptions of crowdsourced context. These categories naturally overlap and are not mutually exclusive, but we highlight them in order to emphasize the different types of knowledge emerging from this body of research. While some papers make multiple contributions and thus span categories, we present them in Table 1 grouped according to their most significant contribution for simplicity.
@@
The most common type of study explored the impact of a CCS on its implementing platform (n=24 papers). These studies all performed analysis of data from X’s Community Notes. These studies asked both causal questions, such as about the effect that displaying a note has on post diffusion (diffusion21; rollout21), as well as descriptive questions about the performance of X’s system in several specific use cases (chujyo25; bouchaudAlgorithmicResolutionCrowdsourced2025b). The studies found strong evidence that CCS can both reduce the spread (diffusion21) and believability (wojcikBirdwatchCrowdWisdom2022b) of misleading posts. The research also argued that the bridging-based algorithm in X’s Community Notes system has demonstrated it can identify high-quality crowdsourced notes that are likely to be perceived as helpful by users across an ideological spectrum (bouchaudAlgorithmicResolutionCrowdsourced2025b). However, the studies also showed that this bridging approach takes a long time to identify “helpful” notes for misleading posts. As a result, there is often a significant delay before notes are displayed to users (renaultCollaborativelyAddingContext2024a; chuaiCommunitybasedFactcheckingReduces2024a; rollout21). Additionally, studies have shown that the vast majority of suggested notes ($\sim 90\%$) are never published due to either a lack of ratings or consensus  (chujyo25; Wirtschafter_Majumder_2023).
@@
Another common type of research study (n=16) examined CCS contributor behavior. While one of these studies used interviews as a data source (yoon25), all of the others relied on the open data from X’s Community Notes. The studies asked descriptive questions about how contributors to the system behaved, such as what types of posts (pilarskiCommunityNotesVs2023; cikm22) and accounts (allen22) contributors tend to leave notes on, what type of content contributors tend to include in notes (jones22; kuuse25), and what types of notes contributors rate as most helpful (solovevReferencesUnbiasedSources2025a; Phillips2025). Three key findings from these studies are that 1) the crowd exerts political bias in the posts that they select to write notes for and the ratings that they assign to notes (allen22); 2) the crowd behaves differently than professional fact-checkers (pilarskiCommunityNotesVs2023), yet often links to them as sources (borenstein2025); and 3) certain types of notes are perceived as more helpful than others, such as those that support claims with links to unbiased (i.e., assessed as politically neutral by a third-party organization [^7] https://mediabiasfactcheck.com/) external sources (solovevReferencesUnbiasedSources2025a).
@@
A third group were design studies that either proposed or built a variation of a CCS (n=10). These papers focused on ways to improve the quantity (liScalingHumanJudgment2025a) and quality (mohammadiAIFeedbackEnhances2025a) of notes, as well as techniques to improve the performance of the note curation algorithm (hawkeye). The bridging-based algorithm for selecting “helpful” notes was introduced in one of these papers (wojcikBirdwatchCrowdWisdom2022b), as a solution to problems of bias (allen22) in a majority-rule note selection system. Recent papers have proposed ways of incorporating AI into the note-writing process and found that LLMs have the potential to generate helpful notes (supernotes; mohammadiAIFeedbackEnhances2025a).
A final group of studies sought to understand user’s perceptions of the notes produced by CCS (n=6). These studies used experiments (liu23), interviews (shusas24), and analysis of X’s data (Hameleers18052024) to answer descriptive questions about how individuals perceived the credibility and helpfulness of context provided by CCS. They found evidence that users perceive notes from CCS as just as, or even more, credible than those from professional fact checkers (liu23; shusas24). While these studies begin to explore the impacts of X’s design decisions, much is still unexplored. As we show next, the potential design space is much wider than what is covered in existing research.
@@ 7.1. Participation
Dimension TikTok Meta Twitter/X YouTube
@@
All users.
Only users from a specific geo/market.
@@
Users.
@@
Users.
Platform employees/contractors.
@@
Nothing: users cannot be context contributors.
Apply.
Receive an invite from the platform.
Be admitted by the platform.
Meet a quality threshold of past performance.
@@
Apply.
Receive an invite from the platform.
Be admitted by the platform.
@@
User-age requirements.
Account-age requirements.
Lack of policy-violations.
Geographic requirements.
Language requirements.
Account authentication requirements.
@@
Yes: platform has discretion over who they admit.
@@
Paid by the platform.
Pseudonymous reputation.
@@
Contributors add notes to posts they organically see on the platform.
Contributors have feeds of posts that need notes, and notes that need ratings.
Platform notifies contributors of specific posts and notes to review.
Users can request notes on posts or additional review on published notes.
@@ 7.2. Inputs
Dimension TikTok Meta Twitter/X YouTube
@@
Free-text inputs
Ability to add source links.
Ability to add references to a portion of the post (e.g. timestamp in the video being noted).
@@
A fixed set of options (i.e., helpful, somewhat helpful, not helpful) to select from.
A free-text explanation field.
A fixed set of explanation options.
@@
All.
Posts by specific account types (public, verified, pages).
Posts of specific types of content (no ads, none that may feature minors).
@@
Must adhere to community guidelines.
Length restrictions.
Cannot be edited once submitted.
@@
Must adhere to community guidelines.
Length restrictions.
Cannot be edited once submitted.
Cannot be deleted once submitted.
@@
Published notes can be hidden by the crowd indefinitely.
Note authors can edit or delete their notes.
@@
Published notes are ”locked” after a certain amount of time.
@@ 7.3. Curation
Dimension TikTok Meta Twitter/X YouTube
@@
Algorithmic decision based on data from contributors.
@@
Platform can remove notes if they violate platform content policy.
@@
Rating contributor’s past platform behavior as a contributor.
@@
Identify notes that appeals across the ideological spectrum (“bridging”).
Identify notes with majority appeal.
Identify factually-accurate notes.
Identify notes that reference reliable sources.
@@
Perhaps the most important design question is “What process determines which notes are shown alongside posts?” The existing systems that have recently grown in popularity all involve algorithmic decisions by bridging-based algorithms that, broadly speaking, identify notes that have been rated as “helpful” by contributors who have disagreed in their past ratings. Even though these bridging algorithms are currently popular, alternatives to this scheme are possible: for example, Twitter’s first implementation of Birdwatch used a majority voting algorithm. Alternatively, it could be possible to achieve bridging goals using contributor information other than their past ratings. Algorithmic approaches are also not the only option, as one could imagine a CCS that selects notes via editorial decisions by either the platform or contributors. Another possible approach is to select notes via digital juries (fanDigitalJuriesCivicsOriented2020) of users, which could be selected either randomly or based on other criteria.
@@ 7.4. Presentation
Dimension TikTok Meta Twitter/X YouTube
@@
Below the text of the original post.
@@
On both the original post and any re-shares of it.
@@
Original post authors.
Contributors who have written/rated the published note.
Every user that replied, liked, or reposted the original post.
@@
At most one.
@@
The author of the note is anonymous/omitted completely.
@@
Only whether or not it is “helpful”.
@@ 7.5. Platform Treatment
Dimension TikTok Meta Twitter/X YouTube
@@
No different from any other posts.
@@
Limit positioning next to ads.
@@
Replaces third-party fact checking program.
Complements platform-provided content moderation.
Notes are presented on posts alongside other “information reliability” indicators.
@@ 7.6. Transparency
Dimension TikTok Meta Twitter/X YouTube
@@
Open source.
Open data.
@@
Not public.
@@
Available to paying customers.
Not public.
@@
Blog posts.
Help articles/FAQs.
Technical documentation.